# GALE: Leveraging Heterogeneous Systems for Efficient Unstructured Mesh Data Analysis

### v1

@@ 4.2.2 Analysis of scalability
Table 3 shows the results obtained when computing critical points. GALE consistently proves to be the fastest data structure, with total computation time halving when the number of CPU threads increases from 8 to 16. However, from 32 threads onward, the time reduction plateaus (more details in Section 4.3.2).
@@
Data structure Consumer number Producer number Initialization time (s) Algorithm time (s) Total time (s) Memory usage (GB)
@@
Similar trends are observed when running the discrete gradient algorithm (see Table 4). GALE delivers the best overall performance, while ACTOPO demonstrates the best scalability. TTK Explicit Triangulation shows minimal speedup, primarily due to limited parallelism during the initialization stage. For GALE, execution time increases significantly with more threads, but it plateaus at 30 consumer threads. The performance gap between GALE and ACTOPO narrows as both share a common initialization routine. In all cases, memory usage remains virtually unchanged, regardless of the number of threads used to run the algorithm.
@@
Data structure Consumer number Producer number Initialization time (s) Algorithm time (s) Total time (s) Memory usage (GB)
@@ 4.3 System Evaluation
In this section, we evaluate the design choices made for GALE, focusing on the time performance throughout the different stages of communication between the CPU and GPU. For the following experiments, we ran all three TTK algorithms and measured the total waiting time of the consumer threads. This waiting time is broken down as shown in Figure 10. For each consumer request, we measure: (1) the time taken to enqueue the request, (2) the time the request spends in the queue before being processed by the leader producer, (3) the data preparation time before launching the GPU kernel, (4) the GPU kernel computation time and the transfer time back to CPU memory, and (5) the time spent integrating data into the data structure for use by the algorithm. For simplicity, we start our analysis with only one consumer thread and will move to multiple consumers later.
@@ 4.3.1 Single consumer
Table 5 presents a detailed breakdown of the waiting time when running the critical points algorithm on the Stent dataset. With only one consumer thread, both pushing a request to the job queue and handling the request are virtually instantaneous. These requests primarily impact Kernel computation and Data integration steps, which are the most time-intensive parts of the pipeline. Notably, the waiting times for the MS-complex computation are significantly higher than the other two algorithms. Overall, this indicates that the queue system is supporting algorithm execution effectively. However, the kernel remains the slowest part, which could be due to the large number of requests or inefficiencies in CPU-GPU communication. The following results will clarify that the problem is not related to CPU-GPU overhead.
@@
Algorithm Critical points Discrete gradient Morse-Smale complex
@@ 4.3.2 Multiple consumers
Table 6 provides a detailed breakdown of the waiting time when extracting critical points from the Stent dataset with different numbers of consumers. Similar trends are observed for the discrete gradient algorithm in Table 7.
@@ 6 Conclusion and Future Works
All supplemental materials are available on OSF at https://osf.io/zxm4w/?view_only=24b44ac1146046c3ada2a3cfcb23626e, released under a CC BY 4.0 license. In particular, they include (1) an Excel file containing the raw data obtained from the experiments and the processed data used for generating figures and tables in the paper, (2) a script to create figures using the processed results and all resulting figure images in PNG format, and (3) a PDF file containing appendices with additional experimental results. The source code of the implementation is also shared in the link above, released under a BSD license.
@@ Acknowledgements.
This work was supported by the National Science Foundation (NSF) award OAC 2403022. This research used in part resources on the Palmetto Cluster at Clemson University under NSF awards MRI 1228312, II NEW 1405767, MRI 1725573, and MRI 2018069. The views expressed in this article do not necessarily represent the views of NSF or the United States government.
@@ A.1 Number of GPU threads for one mesh segment
For the experiment, we extract different types of topological relations for $960$ mesh segments, including $EF$, $ET$, $FT$, $VF$, $VV$, $VE$, and $VT$ relations. We evaluate the performance of different numbers of GPU threads per segment, such as 1, 4, 8, 16, 32, 48, 64, and 96. To reduce the impact of other parameters, we use the largest number of threads in a block ($\leq 1024$) that is a common multiple of all these values, which is 960. To compute the same number of mesh segments, different numbers of GPU blocks will be used.
Since all the experimental datasets show a similar trend, we use the Fish dataset as an example in the paper and place others in supplemental materials. As Figure 11 shows, distributing the computation task to more GPU threads can improve the time performance, but such an advantage is not infinite, i.e., the speedup is hardly noticeable after using more than 32 threads. Since the warp size of the GPU device is 32, using 32 threads ensures that the memory transfer and instruction dispatch of the same segment are grouped into the same warp. Employing more than 32 threads for each segment not only divides the same segment into different warps but also increases the communication and synchronization overhead, thereby limiting the overall speedup. In general, using 32 threads achieves around 3×3$\times$3 × speedup compared to the single-thread execution, except for the $VT$ relation, which can be directly computed from the input tetrahedron list, and more threads need to perform the atomic operation on the same vertex, thereby limiting the speedup.
@@ A.2 GPU block dimension initiated for computation
For the experiment, we have run the selected algorithm with 16 threads and used 8 thread blocks, while varying the number of threads in each block. Since 32 threads have been tested to be the most suitable value to work on one segment in the previous section, the block size is selected as a multiple of 32 (e.g., 32, 64, 128) up to a maximum of 1024 (the limit of our GPU device). Figure 12 shows the time and memory comparison when using different numbers of threads per block. In general, using more threads per block can have more segments precomputed and improve the time performance. However, too many threads may saturate the memory bandwidth and increase the synchronization overhead, which can be easily observed on smaller datasets. Using 512 threads per block achieves relatively the best performance, providing more than 4×4$\times$4 × speedup on average. The difference in memory usage is hard to notice, given that the increased memory for relation arrays represents a relatively small fraction of the total memory.
@@
For the experiment, we have run the selected algorithm with 16 consumer threads and used 512 threads per block, while changing the number of GPU blocks per consumer thread, i.e., 1, 2, 4, 8, 16, and 32. Figure 13 shows the bar charts of time and memory usage. In general, the execution time decreases first and then increases as the number of blocks grows. The main reason is that using more blocks can precompute more mesh segments to reduce the waiting time of consumer threads. However, a higher number of blocks can increase the launch time of the kernel and possibly limit the GPU to run multiple kernels concurrently, taking the leader producer more time to answer the request. The memory usage also increases as the number of blocks grows, however, this is not obvious on larger datasets as an increased number of relation arrays only occupies a small portion compared to the memory space of the mesh. The difference is within 5% on average.
@@ Appendix B Waiting Time Distribution of Multiple Consumers
Figure 14 and Figure 15 show the detailed waiting time distributions for the critical points algorithm using 8, 16, 24, 32, and 40 consumers. The limited variance across consumer threads indicates effective workload balancing. A similar trend can also be observed for the discrete gradient algorithm, as shown in Figure 16 and Figure 17.

### v3

@@ 4.2.2 Analysis of scalability
Table˜3 shows the results obtained when computing critical points. GALE consistently proves to be the fastest data structure, with total computation time halving when the number of CPU threads increases from 8 to 16. However, from 32 threads onward, the time reduction plateaus (more details in Section˜4.3.2).
@@
Data
structure
Consumer
number
Producer
number
Initialization
time (s)
Algorithm
time (s)
Total
time (s)
Memory
usage (GB)
@@
Similar trends are observed when running the discrete gradient algorithm (see Table˜4). GALE delivers the best overall performance, while ACTOPO demonstrates the best scalability. TTK Explicit Triangulation shows minimal speedup, primarily due to limited parallelism during the initialization stage. For GALE, execution time increases significantly with more threads, but it plateaus at 30 consumer threads. The performance gap between GALE and ACTOPO narrows as both share a common initialization routine. In all cases, memory usage remains virtually unchanged, regardless of the number of threads used to run the algorithm.
@@
Data
structure
Consumer
number
Producer
number
Initialization
time (s)
Algorithm
time (s)
Total
time (s)
Memory
usage (GB)
@@ 4.3 System Evaluation
In this section, we evaluate the design choices made for GALE, focusing on the time performance throughout the different stages of communication between the CPU and GPU. For the following experiments, we ran all three TTK algorithms and measured the total waiting time of the consumer threads. This waiting time is broken down as shown in Figure˜10. For each consumer request, we measure: (1) the time taken to enqueue the request, (2) the time the request spends in the queue before being processed by the leader producer, (3) the data preparation time before launching the GPU kernel, (4) the GPU kernel computation time and the transfer time back to CPU memory, and (5) the time spent integrating data into the data structure for use by the algorithm. For simplicity, we start our analysis with only one consumer thread and will move to multiple consumers later.
@@ 4.3.1 Single consumer
Table˜5 presents a detailed breakdown of the waiting time when running the critical points algorithm on the Stent dataset. With only one consumer thread, both pushing a request to the job queue and handling the request are virtually instantaneous. These requests primarily impact Kernel computation and Data integration steps, which are the most time-intensive parts of the pipeline. Notably, the waiting times for the MS-complex computation are significantly higher than the other two algorithms. Overall, this indicates that the queue system is supporting algorithm execution effectively. However, the kernel remains the slowest part, which could be due to the large number of requests or inefficiencies in CPU-GPU communication. The following results will clarify that the problem is not related to CPU-GPU overhead.
@@
Algorithm
Critical
points
Discrete
gradient
Morse-Smale
complex
@@ 4.3.2 Multiple consumers
Table˜6 provides a detailed breakdown of the waiting time when extracting critical points from the Stent dataset with different numbers of consumers. Similar trends are observed for the discrete gradient algorithm in Table˜7.
@@ 6 Conclusion and Future Works
All supplemental materials are available on OSF at https://osf.io/zxm4w/, released under a CC BY 4.0 license. In particular, they include: (1) an Excel file containing both the raw experimental data and the processed data used to generate the figures and tables in the paper; (2) a script for creating figure images from the processed results; and (3) a PDF file with appendices presenting additional experimental results. The source code for the implementation is also provided at the same link, released under a BSD license.
@@ Acknowledgements.
The authors would like to thank General Electric for the Engine dataset, Philips Research (Hamburg, Germany) for the Foot dataset, and Michael Meißner (Viatronix Inc.) for the Stent dataset. The Asteroid dataset was kindly provided by Mathieu Pont [Pont2022wasserstein]. The remaining tetrahedral meshes (Fish and Hole) are courtesy of Yixin Hu from New York University. This work was supported by the National Science Foundation (NSF) under award OAC 2403022. This research used in part resources on the Palmetto Cluster at Clemson University under NSF awards MRI 1228312, II NEW 1405767, MRI 1725573, and MRI 2018069. The views expressed in this article do not necessarily represent the views of NSF or the United States government.
@@ A.1 Number of GPU threads for one mesh segment
For the experiment, we extract different types of topological relations for 960960960 mesh segments, including $EF$, $ET$, $FT$, $VF$, $VV$, $VE$, and $VT$ relations. We evaluate the performance of different numbers of GPU threads per segment, such as 1, 4, 8, 16, 32, 48, 64, and 96. To reduce the impact of other parameters, we use the largest number of threads in a block (1024$\leq$ 1024≤ 1024) that is a common multiple of all these values, which is 960. To compute the same number of mesh segments, different numbers of GPU blocks will be used.
Since all the experimental datasets show a similar trend, we use the Fish dataset as an example in the paper and place others in supplemental materials. As Figure˜11 shows, distributing the computation task to more GPU threads can improve the time performance, but such an advantage is not infinite, i.e., the speedup is hardly noticeable after using more than 32 threads. Since the warp size of the GPU device is 32, using 32 threads ensures that the memory transfer and instruction dispatch of the same segment are grouped into the same warp. Employing more than 32 threads for each segment not only divides the same segment into different warps but also increases the communication and synchronization overhead, thereby limiting the overall speedup. In general, using 32 threads achieves around 3×3$\times$3 × speedup compared to the single-thread execution, except for the $VT$ relation, which can be directly computed from the input tetrahedron list, and more threads need to perform the atomic operation on the same vertex, thereby limiting the speedup.
@@ A.2 GPU block dimension initiated for computation
For the experiment, we have run the selected algorithm with 16 threads and used 8 thread blocks, while varying the number of threads in each block. Since 32 threads have been tested to be the most suitable value to work on one segment in the previous section, the block size is selected as a multiple of 32 (e.g., 32, 64, 128) up to a maximum of 1024 (the limit of our GPU device). Figure˜12 shows the time and memory comparison when using different numbers of threads per block. In general, using more threads per block can have more segments precomputed and improve the time performance. However, too many threads may saturate the memory bandwidth and increase the synchronization overhead, which can be easily observed on smaller datasets. Using 512 threads per block achieves relatively the best performance, providing more than 4×4$\times$4 × speedup on average. The difference in memory usage is hard to notice, given that the increased memory for relation arrays represents a relatively small fraction of the total memory.
@@
For the experiment, we have run the selected algorithm with 16 consumer threads and used 512 threads per block, while changing the number of GPU blocks per consumer thread, i.e., 1, 2, 4, 8, 16, and 32. Figure˜13 shows the bar charts of time and memory usage. In general, the execution time decreases first and then increases as the number of blocks grows. The main reason is that using more blocks can precompute more mesh segments to reduce the waiting time of consumer threads. However, a higher number of blocks can increase the launch time of the kernel and possibly limit the GPU to run multiple kernels concurrently, taking the leader producer more time to answer the request. The memory usage also increases as the number of blocks grows, however, this is not obvious on larger datasets as an increased number of relation arrays only occupies a small portion compared to the memory space of the mesh. The difference is within 5% on average.
@@ Appendix B Waiting Time Distribution of Multiple Consumers
Figure˜14 and Figure˜15 show the detailed waiting time distributions for the critical points algorithm using 8, 16, 24, 32, and 40 consumers. The limited variance across consumer threads indicates effective workload balancing. A similar trend can also be observed for the discrete gradient algorithm, as shown in Figure˜16 and Figure˜17.